OpenStudy (anonymous):

Find the amplitude, period, ohase shift, and endpoint then graph the following function. a) y= cos2(x- pi/2) b) y= 2sin (1/2 x - pi/6)

OpenStudy (luigi0210):

If I'm not mistaken, the formulas are y = Asin(Bx C) + D and y = Acos(Bx C) + D?

OpenStudy (anonymous):

yes for sin and cos

OpenStudy (anonymous):

how do i find the amplitude?

OpenStudy (luigi0210):

The A is amplitude

OpenStudy (anonymous):

so the amplitude is 2?

OpenStudy (luigi0210):

For letter b, yes.

OpenStudy (anonymous):

and now i dont know how to find the period, phase shift, or endpoint.. are you familar with that?

OpenStudy (anonymous):

for letter a, the amplitude is 1?

OpenStudy (luigi0210):

Yup, amp is 1 And it's been years since I've done period, phase, etc

OpenStudy (anonymous):

im using that right now =/ no help

OpenStudy (luigi0210):

maybe @ranga can help you :) I gtg so sorry :(

OpenStudy (ranga):

cos2(x- pi/2) Is the 2 after cos an exponent or just a factor?

OpenStudy (luigi0210):

Good luck!

OpenStudy (anonymous):

its not an exponent

OpenStudy (anonymous):

its a whole number!

OpenStudy (ranga):

ok. take the 2 inside. cos(2x - pi) The period of the function Acos(Bx + C) is 2pi/B So the period of cos(2x - pi) is 2pi/2 = pi Do the same for the sine function.

OpenStudy (anonymous):

but what is the amplitude because there is no number before the cos

OpenStudy (ranga):

A = 1 for cos

OpenStudy (anonymous):

okay for letter b i got 3pi as the period

OpenStudy (anonymous):

can you explain how to find phase shift and endpoint of letter a

OpenStudy (ranga):

period = 2pi/B B = 1/2 for the second function so period = 2p/(1/2) = 2pi * 2 = 4pi

OpenStudy (anonymous):

oh whoops =X how do i find phase shif and endpoint for letter a

OpenStudy (anonymous):

phase shift = c/b so theres no number before x what is b????

OpenStudy (anonymous):

pi/2 ?

OpenStudy (ranga):

For Acos(Bx - C), phase shift is C/B For cos2(x- pi/2) C = pi/2 and B = 1 phase shift = pi/2 do the same for b)

OpenStudy (anonymous):

i got pi/6 / 1/2

OpenStudy (anonymous):

how do i do the math for that???

OpenStudy (ranga):

when you divide by a fraction, multiply by its reciprocal.

OpenStudy (anonymous):

im sorry can you write that out for me?

OpenStudy (anonymous):

pi/6 divided by 1/2 do i write that as p/6 * 2/1

OpenStudy (ranga):

yes. simplify.

OpenStudy (anonymous):

what did you get as answer?

OpenStudy (anonymous):

is it 2pi/6

OpenStudy (ranga):

simplify.

OpenStudy (ranga):

2 is a factor for both top and bottom.

OpenStudy (anonymous):

im sorry i did not understand.. can you show me on the drawing

OpenStudy (ranga):

simplify 2pi/6

OpenStudy (anonymous):

pi/3 ??

OpenStudy (anonymous):

what is endpoint?

OpenStudy (ranga):

To find endpoint for cos2(x- pi/2) solve the inequality: 0 <= 2(x- pi/2) <= 2pi Solve both inequalities and find x.

OpenStudy (anonymous):

can you show me pls

OpenStudy (ranga):

0 <= 2(x- pi/2) <= 2pi divide by 2 0 <= x - pi/2 <= pi add pi/2 pi/2 <= x <= 3/2pi Those are the end points for a) Do the same for b)

OpenStudy (anonymous):

thank you! so muchhhh xx

OpenStudy (anonymous):

do you know how to graph this function?

OpenStudy (ranga):

go to https://www.desmos.com/calculator type in your function as cos2(x- pi/2) on the left column blank space and it will graph it for you.

OpenStudy (anonymous):

thank you :)

OpenStudy (ranga):

you are welcome.